# jBOT: Semantic Jet Representation Clustering Emerges from Self-Distillation

### v1

@@ 4.5 Downstream anomaly detection
Fig. 8 shows the anomaly score distributions, and Fig. 9 shows the ROC curves, where most anomalous signals are reasonably separated from the QCD background. Tab. 4 lists the AUCs for the individual signals and the combined signal. For a baseline comparison, we also quote results from Ref. [43], which uses reconstruction-based autoencoders for this task with different architectures, including a convolutional neural network (CNNAE), a graph neural network (GNNAE), and a Lorentz group equivariant network (LGAE). Our method performance is broadly comparable to the reconstruction-based models and can perform better for some signals. For example, our method using $k$-NN, cosine similarity, and GMM yields AUCs above 0.8 for the W and Z signals, while the reconstruction-based models yield AUCs below 0.8; for the t signal, our method using Mahalanobis distance yields the highest AUC among our metrics at around 0.86, comparing to around 0.89 for CNNAE and GNNAE; for the combined signal, our method using cosine similarity yields AUC of around 0.82 which is close to the highest AUC of 0.83 for LGAE.
It can also be seen that the performance varies with the choice of anomaly score. This is expected because there is a high degree of freedom in defining the metric and hyperparameters for comparing a test example to a large reference set in a high-dimensional embedding space, and different choices probe different aspects of the structure: e.g., local neighborhoods for $k$-NN vs. global distribution for GMM. This flexibility introduces a large space to explore, and optimizations such as hyperparameter scans are required to select the best performing models. Nonetheless, we show that measuring similarity between the test examples and the nominal examples in a self-supervised embedding space pre-trained only on nominal data provides a viable anomaly detection strategy and can yield competitive or even better performance than common reconstruction-based methods.
@@
Table 4: Anomaly detection performance comparing jBOT (particle features only as input) and reconstruction-based autoencoder models from Ref.[43].

### v2

@@ 4.5 Downstream anomaly detection
Fig. 8 shows the anomaly score distributions, and Fig. 9 shows the ROC curves, where most anomalous signals are reasonably separated from the QCD background. Tab. 4 lists the AUCs for the individual signals and for the combined signal. For a baseline comparison, we also quote results from Ref. [43], which uses reconstruction-based autoencoders for this task with different architectures, including a convolutional neural network (CNNAE), a graph neural network (GNNAE), and a Lorentz group equivariant network (LGAE). Our method performance is broadly comparable to the reconstruction-based models and can perform better for some signals. For example, our method using $k$-NN, cosine similarity, and GMM yields AUCs above 0.8 for the W and Z signals, while the reconstruction-based models yield AUCs below 0.8; for the t signal, our method using Mahalanobis distance yields the highest AUC among our metrics at around 0.86, comparing to around 0.89 for CNNAE and GNNAE; for the combined signal, our method using $k$-NN, cosine similarity, and GMM yields AUCs of around 0.82, which is close to the highest AUC of 0.83 for LGAE.
It can also be seen that the performance varies with the choice of anomaly score. This is expected because there is a high degree of freedom in defining the metric and hyperparameters for comparing a test example to a large reference set in a high-dimensional embedding space, and different choices probe different aspects of the structure: e.g., local structure for $k$-NN vs. global structure for GMM. This flexibility introduces a large space to explore, and optimizations such as hyperparameter scans are required to select the best performing models. Nonetheless, we show that measuring similarity between the test examples and the nominal examples in a self-supervised embedding pre-trained only on nominal data provides a viable anomaly detection strategy and can yield competitive or even better performance than common reconstruction-based methods.
@@
Table 4: Anomaly detection performance comparing jBOT (particle features only) and reconstruction-based autoencoder models from Ref.[43].